OpenStudy (anonymous):

Write an equation in point slope form for the line through the given point w/ the given slope (9,-1);m=4/3 Can you please explain how to solve this, but explain it in english? :)

OpenStudy (mathstudent55):

The point-slope form of the equation of a line is: y - y1 = m(x - x1), where m = slope, and (x1, y1) is a point on the line. In this case, the slope, m = 4/3, and the point (x1, y1) = (9, -1), so x1 = 9 and y1 = -1 Take the equation y - y1 = m(x - x1) and just insert the slope and the point (x1, y1) You get y - (-1) = (4/3)(x - 9) y + 1 = (4/3)(x - 9) y + 1 = (4/3)x - 12 y = 4/3 x - 13 There's the equation of the line using the point-slope form, and an explanation in Eglish. I hope it helps.